OpenStudy (anonymous):

Can someone refresh my memory plz?!?!?! What is the Greatest Common Factor of x6 and x4? x x4 x6 x10

jimthompson5910 (jim_thompson5910):

List out the factorization of each x^6 = x*x*x*x*x*x x^4 = x*x*x*x

jimthompson5910 (jim_thompson5910):

Now circle the common factors

OpenStudy (anonymous):

if i do tht there are no common factors its just 4 and 6

jimthompson5910 (jim_thompson5910):

How many 'x' terms are in x^6?

OpenStudy (anonymous):

oh my bad wasn't thinking straight... thnx

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

so the answer is x??

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

x^4

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

since each has at least 4 'x' terms

OpenStudy (anonymous):

yayy!! thnx both of you!!